OpenStudy (anonymous):

I really just need help translating this into an equation... On a buying trip in Los Angeles, Rosaria Perez ordered 120 pieces of jewelry: a number of bracelets at $8 each and a number of necklaces at $9 each. She wrote a check for $1030 to pay for the order. How many bracelets and how many necklaces did Rosaria purchase? 45 bracelets and 75 necklaces 55 bracelets and 65 necklaces 60 bracelets and 60 necklaces 50 bracelets and 70 necklaces

OpenStudy (anonymous):

hold on ill answer it

OpenStudy (anonymous):

i just need help putting it into a formula

OpenStudy (anonymous):

ok ill see to it

OpenStudy (anonymous):

i think it is a simulteanous question

OpenStudy (anonymous):

so i thinnk the formula shuld be: x+y=120 8x+9y+1030

OpenStudy (anonymous):

x+y=120 8x+9y=1030

OpenStudy (anonymous):

Oh. thank you! I kept thinking of it as a mixture problem, not a system, derp

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

thx for the medal